OpenStudy (anonymous):

Solve the given equation by factoring. 2 sin^2 theta + 5 sin(theta) - 12 = 0

OpenStudy (anonymous):

i got \[\pm \sqrt{6}/2 ] and -4

OpenStudy (anonymous):

\[\pm \sqrt{6}/2\] and -4 but what is the theta answers?

OpenStudy (anonymous):

let x = sin θ 2x² + 5x - 12 = 0 Using tic tac toe method to factor, 2x²|-12|-24x² 2x | 4|8x x | -3|-3x (2x - 3)(x + 4) = 0 Zero product property and plug back in. You can take it from there. Look up inverse trig functions if you dunno what to do.

OpenStudy (anonymous):

thetacant be solved for in this equation its basically just "x"

OpenStudy (anonymous):

i need to find all the values adding 2kpi to the answer

OpenStudy (campbell_st):

there are no solutions to this equation as -1< sin < 1